OpenStudy (anonymous):

john is (1)/(3) as old as his father and two years younger than his sister Kathy. The sum of the ages of the three is 67. Find the age of each.

OpenStudy (anonymous):

Set the age of each person relative to each other. Let john's age = x. Let dad's age = y. Let sister's age = z. \[x = (1/3)y\] \[x = z-2\] \[x+y+z = 67\] Substitute: \[x + (3x) + (x+2) = 67\] \[5x +2 = 67\] \[5x = 65\] \[x = 13\] From this we know John is 13, his sister is 15 and his father is 39. Proof: \[13 + 15 + 39 = 67\]

OpenStudy (anonymous):

okaaay. thanks a lot roshan. :)